# An ordering for the strength of functional dependence

### v1

@@ B.8 Proofs of Appendix A
For statement (i), see e.g. [23]; for statements (ii) – (iv), see [4, Lemma A.1]. $\blacksquare$
The equivalence of (i) and (iii) is given in [4, Proposition 2.14]. The equivalence between (ii) and (iii) is a direct consequence of Proposition A.1(ii) and (iii). $\blacksquare$
@@
The equivalence of (i) and (ii) as well as the equivalence of (iii) and (iv) follow from the invariance properties of the convex order in Lemma A.3(iv).
@@
To show that (ii) implies (iii), we consider two cases for $v,v_{n}\in(0,1)$ with $q_{Y}(v)<q_{Y}(v_{n})$ for all $n$. In either case, we obtain for $S_{n}:=P(Y<q_{Y}(v_{n})\mid\mathbb{X})$ and $S:=P(Y\leq q_{Y}(v)\mid\mathbb{X})$ that $\mathbb{E}|S_{n}-S|=\mathbb{E}|P(q_{Y}(v)<Y<q_{Y}(v_{n})\mid\mathbb{X})|=P(q_{Y}(v)<Y<q_{Y}(v_{n}))$. In both cases, we specify $v_{n}$ such that
$\lim_{n\to\infty}\mathbb{E}|S_{n}-S|$ $=P\left(\bigcap_{n\in\mathbb{N}}\{Y<q_{Y}(v_{n})\}\right)-P(Y\leq q_{Y}(v))=0\,.$ (64)
@@
In the first case, let $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})=\mathrm{int}(\overline{\mathsf{Ran}(F_{Y^{\prime}})})$ and $v_{n}\downarrow v$ (i.e., $v_{n}>v$ for all $n$ and $v_{n}\to v$ as $n\to\infty$). Then, $y<q_{Y}(v_{n})$ for all $n$ implies $y\leq q_{Y}(v)$ or $y\in(q_{Y}(v),q_{Y}^{+}(v))$, where $q_{Y}^{+}(t):=\sup\{x\mid F(x)\leq t\}.$ Since $\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ is a $P$-null set for this choice of $v$, we obtain (64).
@@
In the second case, let $v\in(0,1)\setminus\overline{\mathsf{Ran}(F_{Y})}=(0,1)\setminus\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ and $v_{n}\downarrow v_{0}:=F_{Y}(q_{Y}(v))$. Then, we have that $q_{Y}(v)=q_{Y}(v_{0})$. If $F_{Y}$ is strictly increasing on $q_{Y}(v_{0}),q_{Y}(v_{0})+\varepsilon)$ for some $\varepsilon>0$, then $q_{Y}(v_{n})\downarrow q_{Y}(v_{0})$. This implies $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}(v)\geq Y\}$, which yields (64). If $F_{Y}$ is constant on $q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]$ and continuous at $q_{Y}^{+}(v_{0})$, then $q_{Y}(v_{n})\downarrow q_{Y}^{+}(v_{0})$. This implies $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}^{+}(v_{0})\geq Y\}=\{q_{Y}(v_{0})\geq Y\}\cup\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]\}$. Since $\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]\}$ is a $P$-null set, we again obtain the convergence in (64). If $F_{Y}$ is constant on $q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))$ and has a jump at $q_{Y}^{+}(v_{0})$, then $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}^{+}(v_{0})>Y\}=\{q_{Y}(v_{0})\geq Y\}\cup\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))\}$. Once again, using that $\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))\}$ is a $P$-null set, (64) follows.
@@
The convergence in (64) implies $S_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }S$ and $\mathbb{E}|S_{n}|\to\mathbb{E}|S|$. Similar considerations yield $S^{\prime}_{n}:=P(Y^{\prime}<q_{Y^{\prime}}(v_{n})\mid\mathbb{X}^{\prime})\xrightarrow{\penalty 10000\ d\penalty 10000\ }S:=P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ and $\mathbb{E}|S^{\prime}_{n}|\to\mathbb{E}|S^{\prime}|$. Using that $S_{n}\leq_{cx}S_{n}^{\prime}$, we obtain from Lemma A.3(v) that $S\leq_{cx}S^{\prime}$. This proves (iii).
@@
It remains to show that (iii) implies (ii). Therefore, assume that $P(Y\leq q_{Y}(u)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(u)\mid\mathbb{X}^{\prime})$ for all $u\in(0,1)\setminus N$ for some $\lambda$-null set $N\subset(0,1)$. We distinguish between two cases for $v,v_{n}\in(0,1)$ with $q_{Y}(v_{n})<q_{Y}(v)$ for all $n$. Setting $S_{n}:=P(Y\leq q_{Y}(v_{n})\mid\mathbb{X})$ and $S:=P(Y<q_{Y}(v)\mid\mathbb{X})$, we show that
$\lim_{n\to\infty}\mathbb{E}|S-S_{n}|$ $=P(Y<q_{Y}(v))-P\left(\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}\right)=0\,.$ (65)
@@
In the first case, assume that $F_{Y}$ has no jump at $q_{Y}(v)$. Choose $v_{n}\in(0,1)\setminus N$ with $v_{n}\uparrow v$. Then, using left-continuity of $q_{Y}$, we obtain $\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}=\{Y<q_{Y}(v)\}$, which implies (65). In the second case, assume that $F_{Y}$ has a jump at $q_{Y}(v)$ and set $v_{0}:=F_{Y}^{-}(q_{Y}(v))$. The case $v_{0}=0$ is trivial. Hence, suppose $v_{0}>0$. Choose $v_{n}\in(0,1)\setminus N$ with $v_{n}\uparrow v_{0}$. Then, a distinction of cases as in the reverse direction yields $P\left(\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}\right)=P\left(\{Y<q_{Y}(v_{0})\}\right)$, which again yields (65). Both cases imply $S_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }S$ and $\mathbb{E}|S_{n}|\to\mathbb{E}|S|$. Similarly, we obtain $S_{n}^{\prime}:=P(Y^{\prime}\leq q_{Y^{\prime}}(v_{n})\mid\mathbb{X}^{\prime})\xrightarrow{\penalty 10000\ d\penalty 10000\ }P(Y^{\prime}<q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})=:S^{\prime}$ and $\mathbb{E}|S_{n}^{\prime}|\to\mathbb{E}|S^{\prime}|$. Using $S_{n}\leq_{cx}S_{n}^{\prime}$ for all $n$, Lemma A.3(v) yields $S\leq_{cx}S^{\prime}$. This proves (ii). $\blacksquare$

### v2

@@ B.8 Proofs of Appendix A
For statement i, see e.g. [23]; for statements iiv, see [2, Lemma A.1]. $\blacksquare$
The equivalence of i and iii is given in [2, Proposition 2.14]. The equivalence between ii and iii is a direct consequence of Proposition A.1iii and iv. $\blacksquare$
@@
The equivalence of i and ii as well as the equivalence of iii and iv follow from the invariance properties of the convex order in Lemma A.3iv.
@@
To show that ii implies iii, we consider two cases for $v,v_{n}\in(0,1)$ with $q_{Y}(v)<q_{Y}(v_{n})$ for all $n$. In either case, we obtain for $S_{n}:=P(Y<q_{Y}(v_{n})\mid\mathbb{X})$ and $S:=P(Y\leq q_{Y}(v)\mid\mathbb{X})$ that $\mathbb{E}|S_{n}-S|=\mathbb{E}|P(q_{Y}(v)<Y<q_{Y}(v_{n})\mid\mathbb{X})|=P(q_{Y}(v)<Y<q_{Y}(v_{n}))$. In both cases, we specify $v_{n}$ such that
$\lim_{n\to\infty}\mathbb{E}|S_{n}-S|$ $=P\left(\bigcap_{n\in\mathbb{N}}\{Y<q_{Y}(v_{n})\}\right)-P(Y\leq q_{Y}(v))=0\,.$ (B.37)
@@
In the first case, let $v\in\mathrm{int}(\overline{\mathsf{Ran}(F_{Y})})=\mathrm{int}(\overline{\mathsf{Ran}(F_{Y^{\prime}})})$ and $v_{n}\downarrow v$ (i.e., $v_{n}>v$ for all $n$ and $v_{n}\to v$ as $n\to\infty$). Then, $y<q_{Y}(v_{n})$ for all $n$ implies $y\leq q_{Y}(v)$ or $y\in(q_{Y}(v),q_{Y}^{+}(v))$, where $q_{Y}^{+}(t):=\sup\{x\mid F(x)\leq t\}.$ Since $\{Y\in(q_{Y}(v),q_{Y}^{+}(v)]\}$ is a $P$-null set for this choice of $v$, we obtain (B.37).
@@
In the second case, let $v\in(0,1)\setminus\overline{\mathsf{Ran}(F_{Y})}=(0,1)\setminus\overline{\mathsf{Ran}(F_{Y^{\prime}})}$ and $v_{n}\downarrow v_{0}:=F_{Y}(q_{Y}(v))$. Then, we have that $q_{Y}(v)=q_{Y}(v_{0})$. If $F_{Y}$ is strictly increasing on $q_{Y}(v_{0}),q_{Y}(v_{0})+\varepsilon)$ for some $\varepsilon>0$, then $q_{Y}(v_{n})\downarrow q_{Y}(v_{0})$. This implies $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}(v)\geq Y\}$, which yields (B.37). If $F_{Y}$ is constant on $q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]$ and continuous at $q_{Y}^{+}(v_{0})$, then $q_{Y}(v_{n})\downarrow q_{Y}^{+}(v_{0})$. This implies $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}^{+}(v_{0})\geq Y\}=\{q_{Y}(v_{0})\geq Y\}\cup\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]\}$. Since $\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0})]\}$ is a $P$-null set, we again obtain the convergence in (B.37). If $F_{Y}$ is constant on $q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))$ and has a jump at $q_{Y}^{+}(v_{0})$, then $\bigcap_{n\in\mathbb{N}}\{q_{Y}(v_{n})>Y\}=\{q_{Y}^{+}(v_{0})>Y\}=\{q_{Y}(v_{0})\geq Y\}\cup\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))\}$. Once again, using that $\{Y\in(q_{Y}(v_{0}),q_{Y}^{+}(v_{0}))\}$ is a $P$-null set, (B.37) follows.
@@
The convergence in (B.37) implies $S_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }S$ and $\mathbb{E}|S_{n}|\to\mathbb{E}|S|$. Similar considerations yield $S^{\prime}_{n}:=P(Y^{\prime}<q_{Y^{\prime}}(v_{n})\mid\mathbb{X}^{\prime})\xrightarrow{\penalty 10000\ d\penalty 10000\ }S:=P(Y^{\prime}\leq q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})$ and $\mathbb{E}|S^{\prime}_{n}|\to\mathbb{E}|S^{\prime}|$. Using that $S_{n}\leq_{cx}S_{n}^{\prime}$, we obtain from Lemma A.3v that $S\leq_{cx}S^{\prime}$. This proves iii.
@@
It remains to show that iii implies ii. Therefore, assume that $P(Y\leq q_{Y}(u)\mid\mathbb{X})\leq_{cx}P(Y^{\prime}\leq q_{Y^{\prime}}(u)\mid\mathbb{X}^{\prime})$ for all $u\in(0,1)\setminus N$ for some $\lambda$-null set $N\subset(0,1)$. We distinguish between two cases for $v,v_{n}\in(0,1)$ with $q_{Y}(v_{n})<q_{Y}(v)$ for all $n$. Setting $S_{n}:=P(Y\leq q_{Y}(v_{n})\mid\mathbb{X})$ and $S:=P(Y<q_{Y}(v)\mid\mathbb{X})$, we show that
$\lim_{n\to\infty}\mathbb{E}|S-S_{n}|$ $=P(Y<q_{Y}(v))-P\left(\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}\right)=0\,.$ (B.38)
@@
In the first case, assume that $F_{Y}$ has no jump at $q_{Y}(v)$. Choose $v_{n}\in(0,1)\setminus N$ with $v_{n}\uparrow v$. Then, using left-continuity of $q_{Y}$, we obtain $\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}=\{Y<q_{Y}(v)\}$, which implies (B.38). In the second case, assume that $F_{Y}$ has a jump at $q_{Y}(v)$ and set $v_{0}:=F_{Y}^{-}(q_{Y}(v))$. The case $v_{0}=0$ is trivial. Hence, suppose $v_{0}>0$. Choose $v_{n}\in(0,1)\setminus N$ with $v_{n}\uparrow v_{0}$. Then, a distinction of cases as in the reverse direction yields $P\left(\bigcup_{n\in\mathbb{N}}\{Y\leq q_{Y}(v_{n})\}\right)=P\left(\{Y<q_{Y}(v_{0})\}\right)$, which again yields (B.38). Both cases imply $S_{n}\xrightarrow{\penalty 10000\ d\penalty 10000\ }S$ and $\mathbb{E}|S_{n}|\to\mathbb{E}|S|$. Similarly, we obtain $S_{n}^{\prime}:=P(Y^{\prime}\leq q_{Y^{\prime}}(v_{n})\mid\mathbb{X}^{\prime})\xrightarrow{\penalty 10000\ d\penalty 10000\ }P(Y^{\prime}<q_{Y^{\prime}}(v)\mid\mathbb{X}^{\prime})=:S^{\prime}$ and $\mathbb{E}|S_{n}^{\prime}|\to\mathbb{E}|S^{\prime}|$. Using $S_{n}\leq_{cx}S_{n}^{\prime}$ for all $n$, Lemma A.3v yields $S\leq_{cx}S^{\prime}$. This proves ii. $\blacksquare$
Acknowledgement
This research was funded in whole by the Austrian Science Fund (FWF) [10.55776/PAT1669224] project Stochastic orders for functional dependence and [10.55776/P36155] project ReDim: Quantifying dependence via dimension reduction. The second author further acknowledges the support from the WISS 2025 project ‘IDA-lab Salzburg’ 20204-WISS/225/197-2019 and 20102-F1901166-KZP.